Conquest of Elysium

Johan Karlsson Internet: d92jk@efd.lth.se Mail: Magistratsv. 55 X112 226 44 Lund Sweden

March 4, 1996

Abstract

This game is Shareware. Please support it if you want to see more high quality games for the Atari Falcon.

1 installation

You need a colour monitor or tv, an ATARI Falcon and about 3 megs of memory to run this program. If you have got these things the installation will be very simple. If you have a harddrive then all you have to do is copy all files into a single directory and double click on 0start.prg to run the game.

If you don’t have a harddrive or if it is full, you can run this game from floppy. You may already have the game installed on floppy discs, if this is the case you just have to double click on 0start.prg and off you go! If this isn’t the case you will have to copy all the files into the root directory of as many disc as needed. The following files have to be on the same disc: 0start.prg, fant2.bin fntintro.bin mod.dta and all files ending with .wrl.

1.1 troubleshooting

If the display is too large for your tv or rgb monitor you may keep the right mouse button pressed while loading the main part of the game to enter a setup menu. There you may choose pal instead of ntsc to reduce the size of the display.

The intro may have trouble with some VGA monitors. If the intro doesn’t work or you have got bored of it, then just run fant2.prg instead of 0start.prg to skip the intro.

2 Playing the game

Each player controls a powerful warlord or wizard whose objective is to take control over the continent. The game can be played by one to eight players and if there are less than eight players the computer can control some or all of the remaining players.

2.1 Modes of play

After starting the game you will have to choose which game-mode to play. The buttons are Random world, Customized world and Scenario. You may press the buttons by using the mouse or the keys 1–3. How to start the game is very self explanatory, so I won’t write any more about that.

3 Game instructions

When you have selected what kind of game you want to play and what players are participiating you are ready to begin the struggle. First you will see your home citadel in the middle of the screen, surrounded by lots of black unknown terrain. In your citadel you will have one or more commanders, select one by clicking on the square. After that press the button with crossed swords on to select which men are to acompany this commander. Each commander has three action points to spend by walking, spell casting etc. You may walk to any adjacent square not containg water by clicking on it while the commander is selected.

If you press the right button on a square you will see all people that are in that square. If you double click with your right mouse button you will get information on the terrain in that square (or you may single click if there is no army in that square). All locations that are owned by someone will have a coloured shield in the top left corner and all locations containing an army will have a coloured sword in the bottom left corner. There are two other symbols that might appear in a square - there are a cross when the square is blessed and a small temple if a temple has been constructed there.

These are some keyboard shortcuts for commonly used functions.

• q Quit turn

• s Recruit men

• t Change tax

• c Change followers

• p Use special power

• o Use this location’s special power

• 1–4 Change game speed

• s (in fight) speed up fight

• Esc Quit game and return to desktop

• Tab (when using mtos) access desktop

3.1 Tax

The tax determines what your monthly income will be. If you use a large tax for a long period, then the people will eventually revolt. The peoples mood are not only affected by income tax, winter makes them angry, spring makes them happy and rulers using them as sacrifices makes them very angry indeed.

4 Character classes

There are seventeen different charcter classes you may choose from, everyone has different strengths and weaknesses.

4.1 Warlords

4.1.1 Human Warlord

This warlord has a very wide variety of human troops and huge siege engines. He’s the most popular leader among the population, people won’t revolt as quickly against the warlord as against other characters.

4.1.2 Barbarian

The barbarians are physically more powerful than normal men, but they are also more vulerable to magic. The barbarian warlord cannot buy advanced weapons like siege engines.

4.1.3 Maharaja

The Maharaja is an Indian warlord, he usually lives in the south part of the world and uses elephants and tigers in warfare. He may ally with Rakshasas who are powerful tigermen magicians. They require little gold to join but wants elephants to feed on in return.

4.2 Magic Users

4.2.1 Necromancer

This magic user practices the art of creating undead creatures. The Necromancer may raise the dead to join his army, doing this only costs 1 ap per try, but he will become more insane with every attempt to raise the dead. He may also summon the undead by using hands of glory. These hands will be found in larger villages and towns. This summoning of undead is cheaper at some locations e.g temple ruins and old castles. The goal of most necromancers is usually to become a powerful undead and live forever!

4.2.2 Demonologist

The Demonologist sacrifices humans to summon and bind powerful demons. Being a Demonologist is a very dangerous occupation, if the sacrifice is too small the demon will attack the summoner and demons are very powerful indeed.

4.2.3 Warlock

The Warlock uses the forces of the elements to gain his power. For fire he needs rubies, for water he needs saphires, for air he needs diamonds and finally he needs emeralds for the element of earth. By using larger amount of gems he may summon more powerful beings. The most difficult summoning is that of the elemental lords (efreet, marid, djinn and dao) which are extremely powerful commanders that may also summon elemental beings like the warlock himself.

4.2.4 Enchanter

The Enchanter has the ability to animate non-living things. The resources needed for this varies from a silver sword for the creation of a living sword to a complete mine for the creation of a powerful golem. The construction and enchanting of objects require a lot of time and thus costs a lot of action points. The requirements for construction are these.

• necrotod, dead people

• flesh golem, dead people

• tree golem, a forest

• oak golem, an ancient forest

• clay golem, a marsh or swamp

• ice golem, a fresh water lake

• stone golem, a coal mine

• iron golem, an iron mine

• crystal golem, a gold stream

• onyx golem, a silver mine

When constructing a golem the enchanter has to be in the required terrain or next to it if a lake is required.

4.2.5 Witch

The Witch uses fungus to summon and take control of hideous creatures. She may choose the amount of fungus to use in a summoning, the more fungus she uses the easier it will be to control.

4.2.6 Shaman

The Shaman is a master of spirits and mind control. He may use his powers to take control over enemies miles away from himself. If he succeds in taking control over an opponent commander the opponent will retain its commanding skills and special powers. Taking control over a person with high magic resistance is very difficult and requires the Shaman to give him nightmares to prepare him for the possession. When a man has been possessed he will have the advantage of defending if a fight follows.

Spirits may be called upon when brute force is needed to conquer a location. The spirits will disapear after the fighting has been done. Spirits may not appear at uninhabited locations.

4.2.7 Druid

The Druid summons animals from the surroundings to join him. He needs herbs to do these summonings. Which animals are summoned depends on the terrain he summons in, e.g. in the plains there are huge amounts of rabbits and in the jungle tigers may be found. If he has a huge amount of herbs he may attempt to summon legendary beings such as rocs and basilisks in the mountains or ents and beholders in the ancient forests.

4.3 Priests

4.4 Cardinal of El

The priests of El are rewarded by their god depending on how large congregation they have. The religion is very good at healing and blessing, all its priests heal their wounds at a very fast rate and may heal complete armies if the congregation is large enough.

4.4.1 High Priestess of Baal

The High Priestess offers human sacrifices to Baal and in return she gets hideous creatures or powerful cureses on her enemies.

4.5 Non Humans

4.5.1 Troll King

The Troll King is the physically most powerful warlord. He is capable of conquering towns on his own. He may also throw large boulders at city wall just like a catapult. His regeneration enables him to heal at an extraordinary rate, just like his main companions, the trolls.

4.5.2 Orc warlord

The orcs are very strong fighters, however they are also more easily affected by magic than humans. The orc scouts are expert forest burners and may destroy forests to stop people from getting herbs and fungus there.

4.5.3 Dragon Lord

The Dragon Lord may summon dragons to guard his mines. The dragons takes a part of the income in return for their service.

4.5.4 Dwarf King

The dwarves are exelent miners and the Dwarf King get higher income from mines than the other warlords.

4.5.5 Elf Queen

The Elf Queen has the same special abilities as the druid, but they are less powerful.

5 Monsters

There are about three hundred different monsters in this game, so it will probably take a year or two before you have seen them all! Many of the monsters have special abilities (there are over one hundred different special abilities!), here is an exlanation of some of them.

5.1 Spells

Most magic-users are able to cast spells in combat. Spell casting is very difficult if the caster is in meele combat and lesser magic-users won’t be able to cast any spells from the front rank.

5.2 Fear

This magic attack scares one or more opponents if they fail a magic resistance check. If a commander is scared away, he will loose his ability to function as a commander permanently. This also applies to sleep and stun attacks. The Ghost has an extremely powerful fear attack that is impossible to resist, however it is possible to survive it if the victim doesn’t die from fear and his mates win the battle.

5.3 Charm

This is an attack against the opponents magic resistance. Should the attack succeed, the victim will become charmed and join the attacker. A commander that is charmed will loose their ability to command.

5.4 Curse

This is a spell used by some evil religous sects. It turns the victim into a zombie, ready to do it masters bidding. The spell only works on humans, any non-human affected by the spell will simply die.

5.5 Siege engines

Some siege engines may be used to destroy the walls of a fortress without engaging the defenders. When a commander has got one of these machines in his army he gets a catapult button. By pressing this button he may bombard any adjacent square containing a fortress.

5.6 Charge

Knights and cavalry uses a powerful charge to engage the enemy. This charge is only used in the first round of combat. The drawback to this attack is that some spearmen sets has developed a technique to make the charge less favorable. By setting their spears into the ground when the charge come, the knights will charge into these and take some damage.

5.7 Energy drain

Some of the more powerful undeads e.g. Wights and Vampires drain energy from the victim it touches. The energy drained will be used to heal the undead, should it be wounded. A more powerful energy drain called incorporate, takes the victims energy and uses it to grow in strength. By incorporating victims the undead may become stronger than he initially was.

5.8 Immortality

Some powerful beings e.g Vampires and Liches are immortal. When killed in combat they will immediately return to their original home citadel (or their owners citadel if they haven’t got one of their own) and after some rest they will be as new. If the home citadel has been conquered by an opponent, then the immortal will be slayed permanently if it is killed. The goal for many necromancers is to become an immortal being. ”What’s the point of becoming a necromancer, if you can’t live a great many times”.

5.9 How to win

You are out from the game when either all your commanders are killed or when you don’t have a citadel any longer. To win you just have to be the last player left.

6 Credits

This game was created by Magical Science. We are Johan Karlsson and Kristoffer Osterman. To contact us write to the address mentioned at the top of the document. Thanks to Anthony Jacques, Sven Bornemark and Mark Smith for all their help.